OpenStudy (anonymous):

If y=-3 when x=15 , find y when x=-2. PLEASE HELP ME , Write a direct variation equation that relates x and y

OpenStudy (turingtest):

um... there has to be more to this question

OpenStudy (anonymous):

if y=-3 & x=15 then the relation between them is y=x/(-5)

OpenStudy (anonymous):

so!!!!!!! when x=-2 y=2/5

OpenStudy (turingtest):

oh, so we are assuming a direct proportion? good guess Dhashini, bet you're right

OpenStudy (turingtest):

mind you we could make infinitely many functions that give y(15)=-3, so the question as it is posted is incomplete

OpenStudy (anonymous):

LOL

OpenStudy (turingtest):

for instance y=x^2-15x-3 y(15)=-3 y(-2)=31 so this question is incomplete @cashmereonly please post the full question so we can help you to the best of our abilities :)